OpenStudy (anonymous):

How can I solve tough questions related to cube roots of 1 in complex numbers ?

OpenStudy (akashdeepdeb):

Can you please elaborate the question a little?

OpenStudy (anonymous):

In this question (in the attachment) I dont know where to begin

OpenStudy (anonymous):

try writing it as a summation, see if it makes any more sense that way...

OpenStudy (anonymous):

the asker isn't here

OpenStudy (anonymous):

so you want to know how to solve it? ok...

OpenStudy (loser66):

yup

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (n+1)(n+\frac{1}{w})(n+\frac{1}{w^2})\]

OpenStudy (anonymous):

I don't quite understand the part about w and w^2 both being non-real cube roots of 1...

OpenStudy (anonymous):

the nonreal cube roots of 1 are cis(120) and cis(240)

OpenStudy (anonymous):

(cis(120))^2 = cis(240) and (cis(240))^2 = cis(480) = cis(120)

OpenStudy (anonymous):

so... \[\sum_{n=1}^{\infty} (n+1)(n+\frac{1}{-.5+i(.5 \sqrt{3})})(n+\frac{1}{-.5-i(0.5 \sqrt{3})})\] \[\sum_{n=1}^{\infty} (n+1)(n+\frac{1}{-.5+i(.5 \sqrt{3})})(n-\frac{1}{.5+i(0.5 \sqrt{3})})\] simplify this down, then see what falls out?

OpenStudy (loser66):

really? do you really think it should go on that way?OMG

OpenStudy (anonymous):

looks a bit messy, but unless there's a stunning flash of insight I haven't had, that looks like the only way out

OpenStudy (anonymous):

then again, it might be easier to simplify if left in terms of w and w^2...

OpenStudy (anonymous):

ah, yes, that's much better. I plugged (n+1)(n+1/w)(n+1/w^2) into wolfram alpha. it gave me (1+n)(1+wn)(1+w^2n) divided by w^3. I did miss a stunning flash of insight. w^3=1

OpenStudy (anonymous):

still messy, but at least no fractions

OpenStudy (anonymous):

expand (1+n)(1+wn)(1+w^2n)... n^3+(1+w+w^2)n^2+(1+w+w^2)n+1

OpenStudy (anonymous):

@Loser66 can you see or do you know what 1+w+w^2 is equal to?

OpenStudy (anonymous):

the sum of any number of points equally spaced around the unit circle is zero.

OpenStudy (phi):

oops, -1+1= 0

OpenStudy (anonymous):

:) we all make mistakes sometimes

OpenStudy (anonymous):

I have been making one since the beginning of the problem in how I write the summations

OpenStudy (anonymous):

so now we have \[\sum_{k=1}^{n} k^3+1=\sum_{k=1}^{n}k^3+\sum_{k=1}^{n}1\]

OpenStudy (anonymous):

and from here you can see to the end

OpenStudy (loser66):

sorry, friend, above my head, I don't understand a word. hihihi...

OpenStudy (anonymous):

what part don't you understand?

OpenStudy (loser66):

I did as wolframalpha does, make the same denominator, take w out of the sumation and don't know how to do next.

OpenStudy (loser66):

but I have no time to stay. I have class now. Thanks for spending time with me.

OpenStudy (anonymous):

oh. there is a formula to find the sum of the first n cubes. have fun in class

OpenStudy (loser66):

nice to see

OpenStudy (anonymous):

the sum of the first n cubes is (n(n+1)/2)^2 ∑k=1nk3 is essentially the sum of the first n cubes

OpenStudy (anonymous):

@Tushi do you see what I did there?

OpenStudy (anonymous):

Yes and the answer is (in the attachment) Thanks for helping me. You have come pretty close to the answer (just a +4n term is missing) How can we get the answer ?

OpenStudy (anonymous):

simplify (n(n+1)/2)^2 +n

OpenStudy (anonymous):

do you see it?

OpenStudy (anonymous):

Yes ... Thanks!!! But the formula for sum of first n cube roots is (n(n+1)/2)^2, where did you get the extra n from ?

OpenStudy (anonymous):

\[\sum_{k=1}^{n} k^3+1=\sum_{k=1}^{n}k^3+\sum_{k=1}^{n}1\]

OpenStudy (anonymous):

\[\sum_{k=1}^{n}1\] equals n

OpenStudy (anonymous):

Now I understand. Thanks a lot!

OpenStudy (anonymous):

you're welcome this was actually a fun problem...

OpenStudy (anonymous):

I find these summation problems quite difficult and confusing ... Can you help me a bit? You seem to be a pro in summation.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What are the basic formulas for summation? Our school textbooks assume we know all summation formulae without giving them anywhere in the book and use them frequently. I always get lost after the sigma sign.

OpenStudy (anonymous):

Thank you... Do I have memorize all of them ?

OpenStudy (anonymous):

you'll get it instinctively with time and practice

OpenStudy (phi):

If you have to answer questions like this one without reference to books , it would appear so. on the other hand, if you have access to google, then no.

OpenStudy (anonymous):

do you want to practice a bit here?

OpenStudy (anonymous):

That would be nice !! Where can I get the questions from ?

OpenStudy (anonymous):

i'll make them up

OpenStudy (anonymous):

\[\sum_{1}^{7}2n+1\]

OpenStudy (anonymous):

Now i am lost

OpenStudy (anonymous):

ok, first try to split it up into two sums of just one term

OpenStudy (anonymous):

7 Summation 1 2n + 7 Summation 1 1 Is that correct ? Where is the summation button on the keyboard ?

OpenStudy (anonymous):

oh, it's in the equation button below vvv down there

OpenStudy (anonymous):

yes, you split it up correctly

OpenStudy (anonymous):

so now, do you know what to do with the sum from 1 to 7 of 1?

OpenStudy (anonymous):

Do I add 1+1+1+1... 7 times to get 7?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

now what do you do with the other one?

OpenStudy (anonymous):

No idea :(

OpenStudy (anonymous):

it's simple enough you could add it up manually or you could move the 2 outside the summation \[\sum_{1}^{7}2n=2\sum_{1}^{7}n\]

OpenStudy (anonymous):

do you know the formula for the sum of the first n numbers?

OpenStudy (anonymous):

n(n+1)/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you see why you can move the constant term (2) outside the summation?

OpenStudy (anonymous):

and can you find the final answer now?

OpenStudy (anonymous):

How can we move 2 outside the summation? Dont we have to add it 7 times?

OpenStudy (anonymous):

the sum from 1 to 7 of 2n is 2+4+6+8...14 we can factor this... 2(1+2+3...7)

OpenStudy (anonymous):

we can take it out because it's always the same

OpenStudy (anonymous):

sorry, it's my bedtime now. I live in China.

OpenStudy (anonymous):

Oh sorry ... I forgot my basic algebra. So we get the answer 2 x n(n+1)/2 = n(n+1) We put n=1, add it to n=2, add it to n=3, ... add it to n=7. Will we get the answer then?

OpenStudy (anonymous):

when you replaced the summation with the (n(n+1))/2 you now only have to put in the highest n

OpenStudy (anonymous):

so just put in n=7 and don't forget the 7 we had earlier

OpenStudy (anonymous):

goodnight

OpenStudy (anonymous):

And we get 7 x 8 = 56. Thanks for helping me and goodnight!

OpenStudy (anonymous):

you got this

OpenStudy (anonymous):

(just don't forget the 7 we had earlier...)

OpenStudy (anonymous):

What do we do with that 7? just put it for n, isn't it?